If the marginal cost to make a good is $20 and the price elasticity of demand is -6 what price should be charged via the optimal markup rule?

Enter as a value.
Answer #1

The formula for calculating the optimal markup pricing is

= P=MC*(E/1+E)

E = -6

MC = 20

P = 20*(-6/1-6)

= 20*6/5

= 24